# DynaGuide: Steering Diffusion Polices with Active Dynamic Guidance

### v1

@@ A.1 Steering in a Simple 2D Block Task
A.2 Ablations and Hyperparameter Investigation
@@
In practice, we use a guidance strength and $\sigma$ chosen from a gridsearch for each experiment, although the result of the gridsearch generally yields $s\in[1.2,1.8]$ and $\sigma\in[20,40]$. On a new task, it is sufficient to pick starting values of $s,\sigma$ from these ranges and adjusting them based on performance. We use stochastic sampling $M=4$ for our experiments as a balance of stability and computation efficiency. We use $20$ guidance conditions per task, which is more than sufficient for Switch-On but is more important for underspecified goals (§4.2) and harder objectives.
@@ B.1 DynaGuide Implementation Details
Inference-Time Steering. Before inference-time, we take the guidance conditions and pre-compute their embeddings. Then, during inference-time, we compute the objective (Eq. 2). Empirically, the Euclidean distance yields more stable results in lieu of squared distance in Eq. 2. We backpropagate this metric through the dynamics model and use the gradient with respect to the input actions as the guidance signal. We find $\bar{\alpha}$ using the DDIM scheduler. To implement stochastic sampling, we repeat the same denoising step multiple times within the denoising loop.
@@
For more details on the exact implementation, refer to the submitted code.
@@ B.3 Distance Metric: Derivation and Design Choices
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d=\sum-\frac{1}{2\sigma}||z_{i}^{+}-\hat{z}_{t+H}||_{2}^{2}}$
@@
$=\log\sum_{i}\exp\left(-\frac{1}{2\sigma}||z_{i}^{+}-\hat{z}_{t+H}||_{2}^{2}+C\right)$
$=\log\sum_{i}\exp(C)\exp\left(-\frac{1}{2\sigma}||z_{i}^{+}-\hat{z}_{t+H}||_{2}^{2}\right)$
$=C+\log\sum_{i}\exp\left(-\frac{1}{2\sigma}||z_{i}^{+}-\hat{z}_{t+H}||_{2}^{2}\right)$
@@
$\mathbf{d}=\log\left[\sum_{i}\exp\frac{-||\phi(g_{i}^{+})-h_{\theta}(\phi(o_{t}),\mathbf{a})||^{2}_{2}}{\sigma}\right]-\log\left[\sum_{j}\exp\frac{-||\phi(g_{j}^{-})-h_{\theta}(\phi(o_{t}),\mathbf{a})||^{2}_{2}}{\sigma}\right]$

### v2

@@ A.1 Steering in a Simple 2D Block Task
Red Grey
DynaGuide 75% 25%
DynaGuide-Sampling 60% 40%
Base Policy 55% 45%
Table 2: CupPreference Additional Result We add a baseline to the GuideRed condition of CupPreference (Section 4.5. While DynaGuide-Sampling improves over baseline, DynaGuide achieves the best guidance.
@@
A.2 Additional Goal-Conditioning Baselines
In Experiments 1-3, we use a simple goal conditioning baseline by using the last trajectory state as the goal representation. To maintain fairness with the base policy, we also train the goal conditioned model on only the test environment’s data in Calvin. However, we note that the dynamics model in DynaGuide is trained with multi-environment data. To maintain fairness with the total data exposed to all parts of a method, we train another goal-conditioned policy on the same multi-environment data. When evaluated on UnderspecifiedObjectives, this additional data improved performance across some steering targets, although DynaGuide still outperformed this augmented goal conditioning baseline (GC w/ All Data, Table 3).
Finally, we try to apply an algorithmic improvement on the goal conditioning baseline by implementing Hindsight Experience Replay [2]. Instead of conditioning the policy on only end states during training, we also sample intermediate future states to broaden the goal distribution. We find that this improves steering success across some tasks but not others (GC w/ HER, Table 3). Overall, we find that goal conditioning can be improved for select tasks with additional data or algorithmic changes, but these augmentations do not raise the performance of this baseline to that of DynaGuide.
button_on button_off switch_on switch_off drawer_open drawer_close door_left door_right
[…]
Table 3: Stronger Goal Conditioned Methods. Strengthening the goal conditioning baselines with additional data or hindsight replay [2] shows some improvement in steering ability, but as tested on UnderspecifiedObjectives (Exp 3), they still perform worse than DynaGuide.
A.3 Additional Real Robot Baseline
The real robot experiments (Section 4.5) demonstrated the ability for DynaGuide to steer pretrained policies. To further demonstrate the performance of DynaGuide on real robots, we add DynaGuide-Sampling baseline for one real robot task. As seen in Table 2, influencing the action sampling with the diffusion model leads to an improved success rate of the target task over baseline, but the classifier guidance of DynaGuide still yields a higher success rate of the target task.
A.4 Ablations and Hyperparameter Investigation
@@
Task Scale $\sigma$
switch_on 1.5 30
switch_off 1.5 30
drawer_open 1.0 40
drawer_close 1.0 40
button_on 1.0 30
button_off 1.0 30
door_left 1.5 30
door_right 2.0 15
Table 4: Sampling Hyperparameters per Task. Per environment, the optimal parameters are very similar across tasks.
In practice, we use a guidance strength and $\sigma$ chosen from a gridsearch for each experiment. These parameters are often very similar across tasks in the same environment (Table 4), which means that switching target tasks will not require extensive hyperparameter search. We use stochastic sampling $M=4$ for our experiments as a balance of stability and computation efficiency. We use $20$ guidance conditions per task, which is more than sufficient for Switch-On but is more important for underspecified goals (§4.2) and harder objectives.
@@
In experiments 1-5, we leverage a more diverse dataset collected across four different Calvin environments to train the dynamics model. This data magnitude improves training stability and reduces overfitting. However, because we test only on a single Calvin environment, we conduct a data ablation and train the dynamics model only on data from that environment. As seen in Table 5, reducing the dynamics training data also reduces the steering success, although the performance is still above baseline (Experiment 4.1). This ablation demonstrates that DynaGuide can transfer information from related environments to help steer a policy in a target environment, thereby reducing the data requirement for training a dynamics model in a new environment.
button_on button_off switch_on switch_off drawer_open drawer_close door_left door_right
[…]
Table 5: Reducing Training Data in Dynamics Model. Ablation result from ArticulatedParts (Section 4.1) using DynaGuide with a dynamics model trained only on the D-split of Calvin.
A.5 Classifier-Free Guidance Baseline
An additional relevant baseline is the Classifier Free Guidance (CFG) [20] that leverages versions of the diffusion policy itself to guide the diffusion process. We implement and run CFG as a baseline for MultiObjectives. CFG can guide towards multiple objectives, a feat that the goal-conditioned baseline was not able to accomplish. Table 6 presents these additional results for a CFG baseline, which should be compared to those of DynaGuide and DynaGuide-Sampling in Figure 5. As expected, CFG can successfully guide the policy towards multiple objectives, including accomplishing multiple behaviors and avoiding other behaviors. However, the total success rate for positive guidance (first three rows in Table 6) is lower than that of DynaGuide (Figure 5).
The CFG is also able to steer the policy away from behaviors, as seen in the last three rows of Table 6. The success rates of the undesired behaviors for CFG and DynaGuide are within one standard deviation. However, while DynaGuide nearly always offers a valid alternative behavior (Figure 5), the CFG baseline fails to execute a behavior more than 10% of the time for all three tested conditions (No Behavior, Table 6). This is an expected result, as the log-sum-exp latent selection method of DynaGuide allows the guidance to find an appropriate behavior and minimize the influence of other contradictory behaviors. In contrast, the CFG implementation gives equal weight to all guidance signals. The lack of a latent distance makes it difficult to select the most relevant guidance. Contradictory guidance signals can influence the diffusion process negatively, leading to poorly-formed actions and unsuccessful behaviors.
Button Switch Drawer Door Blocks No Behavior
[…]
Table 6: Additional Classifier-Free Guidance (CFG) Baseline. Behavior distribution for a CFG baseline on the MultiObjectives Experiment shows that CFG can guide towards multiple objectives but struggles with avoiding specific behaviors without an increase in overall failures (lower three rows).
@@ B.1 DynaGuide Implementation Details
Inference-Time Steering. Before inference-time, we take the guidance conditions and pre-compute their embeddings. Then, during inference-time, we compute the objective (Eq. 2). Empirically, the Euclidean distance yields more stable results in lieu of squared distance in Eq. 2. We backpropagate this metric through the dynamics model and use the gradient with respect to the input actions as the guidance signal. Like the base policy, we use the DDIM sampler to reduce inference-time computation from 100 steps to 10 steps. To implement stochastic sampling, we repeat the same denoising step multiple times within the denoising loop.
@@ B.2 Implementation of Baselines
Classifier-Free Guidance [20]. We use the same goal-conditioned architecture as the goal-conditioned baseline. However, during training, we randomly zero the goal condition to represent the unconditional input needed in classifier-free guidance. During inference, we compute the classifier-free guidance noise estimation using the same conditions used in the goal-conditioned policy baseline.
@@ B.3 Distance Metric: Derivation and Design Choices
[…]
@@
[…]
@@
[…]
@@
In practice, $\mathbf{g}^{+}$ can be either a future state $(t+H)$ or a final state ($T$). Using a future state $t+H$ is easier to train for the dynamics model but it is harder to find a correct $\mathbf{g}^{+}$ during inference-time guidance, as it must be a state achievable $H$ steps from the current timestep. Alternatively, predicting a final state $T$ means that final states from demonstration trajectories can be used as $\mathbf{g}^{+}$, making the choice of inference-time guidance examples straightforward. However, training a dynamics model to predict final state can be difficult for long-horizon tasks. We find that the Calvin environment enables end state prediction ($T$), but the real world robot environment requires setting $H=48$ steps into the future. We feed entire trajectories as $\mathbf{g}^{+}$ and rely on the log-sum-exp latent selection to dynamically pick the most relevant guidance points. The choice of $\mathbf{g}^{+}$ is an important design choice and it depends on the environment and the data.